IODP Proceedings    Volume contents     Search



The Chikyu left Hachinohe Port at 1200 h on 26 July 2012 and arrived at IODP Site C0020 after 7.5 h of transit. The corrosion cap of Hole C9001D, which was drilled to 647 m drilling depth below seafloor (DSF) and cased to 511 m DSF during the Chikyu shakedown cruise in 2006, was retrieved at the surface by 2400 h on 27 July. Ten transponders were installed on the seafloor by the remotely operated vehicle (ROV) on 28 July, and failure mode effect analysis for the Acoustic Position Reference System was completed on 29 July.

While conducting a surface pressure test of the blowout preventer (BOP), the science party was shuttled aboard by helicopter flights on 31 July. Technical problems ensued during the BOP function test, and troubleshooting continued until the function test was completed at 1330 h on 6 August.

After a successful pressure test, the BOP was transferred to the well center at 0130 h on 7 August and was lowered to the moonpool. However, the BOP control system did not function correctly, so the BOP was picked up on the cart. Incorrect connection of the control line conduit manifold was rectified on 8 August. Running of the BOP and riser started at 1300 h on 8 August.

BOP landing was attempted at ~2130 h on 11 August, but the ROV image was lost just before the moment of landing. A successful landing was confirmed before 2200 h when the ROV image was recovered.

Extensive BOP tests were carried out after the landing until completion of the pressure test on 14 August. Cement was drilled out with a 17½ inch bottom-hole assembly (BHA), and seawater was evacuated starting at 1515 h on 14 August and reached 1854.5 m drilling depth below rig floor (DRF), 1 m above the bottom of the cement. Calcium carbide was added to the drilling mud in order to check the annulus volume. Old mud in the hole and riser was displaced with 1.06 sg KCl-NaCl/polymer/PPG (KNPP) mud. A formation integrity test at the 20 inch casing shoe (1719.5 m DRF) was carried out.

Drilling into fresh formation started on 15 August. Drilling a 17½ inch hole from 1854 to 2171 m DRF (645.5–962.5 m DSF, 317 m) was achieved at an average rate of penetration (ROP) of 1.23 m/min. A maximum mud-gas concentration of 18.94% was observed at 1991 m DRF. After circulation and bottoms up, drilling resumed from 2171 to 2208.5 m DRF (962.5–1000 m DSF, 37.5 m) at 0115 h on 16 August. The average ROP was 1.38 min/m. A high mud-gas concentration (>10%) was occasionally observed. Drilling from 2208.5 to 2375 m DRF (1000–1166.5 m DSF, 166.5 m) proceeded at an average ROP of 1.45 min/m.

At 2200 h on 16 August, we encountered total mud loss at 2375 m DRF. The mud loss rate decreased with time to 8.5 m3/h at 2400 h. Lost circulation material (LCM) was spotted (12 m3) at 0230 h on 17 August. The mud loss rate decreased to 0.8 m3/h after the second spotting of LCM at 1200 h but continued through the day. A decision was made to continue drilling with seawater gel mud.

Drilling resumed at 0515 h on 18 August from 2375 to 2400 m DRF (1166.5–1191.5 m DSF, 25 m). The average ROP was 1.78 min/m. Mud return was confirmed and the pump rate was adjusted. Drilling continued to 2453.5 m DRF (1245 m DSF), and then to 2471.5 m DRF (1263 m DSF) with a sweep of hi-vis mud and circulation bottoms up.

The drilling BHA was pulled to the surface at 1500 h on 19 August. The nominal seat protector was retrieved at 2100 h, and preparation of 13⅜ inch casing began. Running 13⅜ inch casing started at 0400 h on 20 August, and running of the casing hanger with running strings started at 1630 h. However, the casing hanger was pulled back to the surface because it was improperly made up. The casing hanger landed on the wellhead at 0015 h on 21 August. Cementing of the casing began at 0515 h. The pressure test of the BOP was completed on a second attempt, after the seal-retrieving tool was set.

A 12¼ inch drill-out cement BHA was made up at 2030 h on 22 August and run into the hole to 2384 m DRF. Casing pressure tests were carried out twice, and cement was drilled out to reach 2.5 m into new formation at 1830 h on 23 August. A leak-off test was completed after drilling fluid was displaced from seawater gel to KNPP mud. Preparation for a 10⅝ inch rotary core barrel (RCB) BHA was started at 1600 h on 24 August and running into the hole to 2444 m DRF was finished at 0700 h on 25 August. After circulation and bottoms up, the center bit was dropped and a 10⅝ inch hole was drilled 11 m into the formation to reach 2485 m DRF (1278.5 m DSF).

Coring operations started at 1600 h. Core 337-C0020A-1R was recovered on deck at 1822 h, with 2.23 m of recovery from 9.5 m of drilling advance. The following core (2R) was recovered at 2123 h, with improved recovery of 4.08 m. Core 3R was cut after drilling to 2578.5 m DRF (1370 m DSF). A short advance of 5.0 m was attempted to improve core recovery and quality. A good 3.68 m core was recovered at 0929 h on 26 August. Core 4R recovered 3.75 m from 9.5 m of advance. After reviewing the core recovery and quality, short advances were employed thereafter. Core 5R was cut after drilling to 2698.5 m DRF (1490 m DSF). The 2.7 m long core was on deck at 0504 h on 27 August. The following core (6R) was a successful recovery of 4.12 m from 5.0 m advance. After Core 6R, drilling was intended to reach 2841.5 m DRF (1633.0 m DSF), but before reaching the target depth, a low-ROP interval was encountered, and the decision was made to core to determine the material of the hard formation. Core 7R was cut from 2807.5 to 2812.5 m DRF (1599.0–1604.0 m DSF) and 1.15 m was recovered at 1915 h. The core included gravel of volcanic origin and showed a different lithology from previous cores. The Co-Chief Scientists decided to take a large-diameter coring (LDC) system core at this depth, targeting the presumed Oligocene/Eocene boundary that was observed in the low-ROP interval in a nearby well.

The RCB BHA was pulled to the surface at 0830 h on 28 August. Running the Baker Hughes INTEQ 8½ inch industrial coring BHA started at 1500 h. LDC cutting started at 1045 h on 29 August. Coring operations were stopped at 2834 m DRF at 1530 h (before reaching 27 m of drilling advance in the original plan) because core jamming was suspected when an increase in pump pressure and no penetration were observed. The LDC core was recovered on deck at 0730 h on 30 August. Core 8L recovered 10.0 m from 21.5 m of advance. The core was cut into 1.0 m long sections at the middle pipe rack and transferred to the laboratory.

As LDC did not reach the target depth, it was decided to take the next core at 2834 m DRF. Running a 10⅝ inch RCB BHA started at 1115 h. The LDC interval was opened from 8½ inches to 10⅝ inches, and RCB coring resumed at 0545 h on 31 August. Slow penetration continued, and drilling advance was stopped at 4 m for Core 9R, which was 3.95 m long and was recovered on deck at 0813 h. Core 10R was on deck at 1140 h, and 5.41 m was recovered from 9.5 m of advance. Based on a preliminary inspection of the sample, no particular change in lithology and age was inferred. The Co-Chief Scientists decided to stop coring and drill instead to 1760 m DSF.

Before reaching the next coring interval, an increase in ROP was observed, suggesting a change in lithology. Three consecutive RCB cores were cut from 1737.5 m DSF with full 9.5 m of advance. Cores 11R–13R were recovered on deck on 1 September and showed good recovery. Core 14R was cut after drilling to 1820 m DSF and was recovered at 1843 h. Again, there was not a notable change in lithology. The Co-Chief Scientists decided to drill to 1919 m DSF, at which depth the coal-bearing interval was considered to start in the revised interpretation of seismic profiles based on the observed drilling parameters and cuttings.

Four consecutive RCB cores were cut from 1919 m DSF. Core 15R was on deck at 1230 h on 2 September and contained a thick coal layer. The following Cores 16R (with advance of 7.5 m) and 17R (with full advance) were mainly loose sand. During cutting of Core 18R, the drill bit was suspected to be worn out, and the core was recovered on deck at 2153 h after a drilling advance of 4.5 m. It was decided to pull out of the hole, and the drill bit was retrieved on the surface at 1330 h on 3 September. Three nozzles were plugged in the drill bit.

After Core 18R, it was also decided that RCB coring would continue and LDC would be canceled. The reasons for the change in plans were that (1) RCB coring had shown unexpectedly high core recovery in coal-bearing lithology, at least as high as LDC, (2) core quality of coal recovered by the RCB was good enough to fulfill our scientific objectives, (3) uncertainties in LDC operations risked jeopardizing the expedition goals in the limited time remaining for the site operations, and (4) quicker RCB operations allowed for a longer total coring interval and more flexibility in the future analyses of the cores.

A new drill bit was installed and run into the hole starting at 1430 h. Coring operations resumed at 1000 h on 4 September and continued smoothly to take seven consecutive RCB cores from 3158.5 to 3212.0 m DRF (1950–2003.5 m DSF). Core recovery was high, and coal-bearing sequences were obtained. As various lithologies within and around the coal-rich formation were available for sampling, the Co-Chief Scientists concluded that we had fulfilled our operational mission in this interval. They decided to drill deeper to investigate the broader range of the hydrocarbon system and to explore the limits of life.

Drilling with a center bit from 2003.5 m DSF started at 0715 h on 5 September. Core 26R was cut from 3318.5 to 3328 m DRF (2110–2119.5 m DSF; 9.5 m). At 2111 m DSF, Hole C0020A became the deepest hole in scientific ocean drilling. On 6 September, we celebrated drilling a record-breaking core with >100% recovery of 9.58 m. Spot coring continued at 100 m intervals, and Cores 27R, 28R, and 29R were taken from 2200, 2300, and 2400 m DSF, respectively.

The final total depth (TD) was anticipated at 2460 m DSF, and the Co-Chief Scientists requested to take the last pair of cores with 9.5 m and 4.0 m advances. While cutting the last core (31R), jamming was suspected at 0.5 m of advance. The Operations Superintendent (OSI) decided to retrieve the core at that point. While waiting for recovery of the short core, scientists observed unique features in the X-ray computed tomography (CT) image of Core 30R. The Co-Chief Scientists and OSI decided to take another full core from 2456.5 m DSF. At 0354 h on 9 September, the last core (32R) was on deck. TD of Hole C0020A was 2466 m DSF.

After pulling out of the hole, wireline logging operations started. Logging Run 1 (Platform Express–High-Resolution Laterolog Array–Hostile Environment Natural Gamma Ray Sonde) started at 0200 h on 10 September. Data quality was very good, and the caliper showed that the hole condition was good through the entire interval. Logging Runs 2 (Formation MicroImager–Dipole Sonic Imager–Environmental Measurement Sonde [EMS]–gamma ray [GR]) and 3 (combinable magnetic resonance-GR) followed. High-permeability layers were selected for use of the Modular Formation Dynamics Tester (MDT) based on the results from the first three runs. After a wiper trip, logging Run 4 (MDT-GR) started at 2230 h on 12 September. Pretests for fluid mobility and formation pressure measurements were carried out at 31 horizons. Formation fluid samples were taken at six horizons of high mobility, and the six bottles were recovered on deck at 0510 h on 14 September. The sample bottles were delivered to the laboratory during logging Run 5 (vertical seismic profile [VSP]-GR). The last run was finished at 1815 h, ending scientific operations on the rig floor.

This expedition was originally scheduled for March–May 2011 but was postponed because of the Tohoku-oki earthquake and the following tsunami hazard that hit the eastern coast of Japan, including Hachinohe Port. The Chikyu suffered damage on the ship body and lost one of the thrusters during emergency evacuation from the port. Repair and reinstallation of the thruster were completed in June 2012, and this expedition was rescheduled for implementation in July 2012.